# Leveraging LLM-based agents for social science research: insights from citation network simulations

### v1

@@ 1 Main
Figure 1: An Illustration of One Simulation Step in the CiteAgent Framework. a, Initialization: LLM-based agents are built as distinct authors, each endowed with distinct attributes, denoted as AA; b, Socialization: We designate an active author subset, denoted as $A_{a}$. Each active agent $a\in A_{a}$ engages in a group discussion with collaborators and collaboratively develops paper drafts. c, Creation: Each active agent utilizes a scholarly search engine to retrieve relevant papers and finalize the paper drafts with reference selection.
@@
CiteAgent includes two entity types: the author set $A={a_{1},\dots,a_{i},\dots a_{n}}$ and the paper set $P={p_{1},\dots,p_{j},\dots p_{m}}$. We employ LLM-based agents to initiate authors and mimic human behaviors. We model interactions between these entities to trace the evolution of citation networks. We initialize the author set AA and paper set $P$ from the citation network dataset $G_{0}$, which is enriched with seven paper attributes (e.g., paper content (abstract), paper timeliness, author name, author country, paper citation, paper topic, and author citation) and five author attributes (e.g., citation, institution, expertise, topic, and nationality). Following the $G_{0}$ setup, we run iterative simulation steps in CiteAgent, as shown in Figure 1. Each step integrates new papers and authors into AA and $P$ following three main stages: Initialization, Socialization, and Creation. In the initialization stage, we build LLM-based agents to instantiate authors. We update AA with new authors, and adjust for overloaded authors who have published excessively. Moving on to the socialization stage, authors collaborate, sharing insights and develop paper drafts. Finally, in the creation stage, authors finalize drafts by referencing relevant papers and updating $P$ with completed papers. As simulation steps proceed, $G_{0}$ progressively expands into a larger network. These steps represent the evolutionary process of the citation network. For validating the simulation authenticity of CiteAgent, we verify that the generated citation network conforms to several classic network science theories, including the power-law distribution [1], citational distortion [22] and shrinking diameter [28] phenomenon.
@@ 2.4 Reasons Behind Citational Distortion
To address this issue, we propose a new metric that normalizes for the cumulative advantage of paper volume. Inspired by the Institutional Preference Score (IPS), which was introduced to capture the citation preferences of various institutions for AI papers in the industry [19]. Likewise, we propose the Referencing Preference Score (RPS) to examine the referencing preferences of agent set AA for papers from country set $C$:
@@
Here, $\text{RPS }(C)>1$ demonstrates a preference for papers from $C$ than would be expected under random referencing behaviors. If reference selection is determined by preferential attachment based on paper counts, we would expect the referencing probability to mirror the distribution of these counts, resulting in an RPS of 1 for all countries (if the ref. share of $C$ in papers is zero, we default to $\text{RPS}(C)=1$. We repeat all simulation experiments three times to calculate the RPS standard deviation error. As shown in Figure 5c, for peripheral countries, we find that $\text{RPS}(C_{core})$ is comparable to, or even slightly lower than, $\text{RPS }(C_{peripheral})$. Moreover, as illustrated in Figure 5d, in the simulations of GPT-4o-mini and LLAMA-3-70B based authors, $\text{RPS}(C_{peripheral})$ gradually exceeds $\text{RPS}(C_{core})$ over the years. In contrast, in the simulations of GPT-3.5 based authors, $\text{RPS}(C_{\text{peripheral}})$ consistently remains slightly lower than $\text{RPS}(C_{\text{core}})$. Thus, in the CiteAgent simulations, authors exhibit no significant preference for papers from peripheral countries over those from core countries; in fact, they show a slight inclination toward papers from peripheral countries. This observation is consistent with the citation motivation statistics presented in Figure 4a and the results of anonymous-public counterfactual experiments, further confirming the absence of preference for papers from core countries over those from peripheral countries. Therefore, the observed inequality in international citations primarily stems from the uneven distribution of researchers across countries, particularly the concentration of authors in core countries, rather than from intentional biases based on the authors’ nationality during the reference selection.
@@ 4 Methods
Before simulation steps begin, CiteAgent undertakes three configuration setups: (1) Time Flow: CiteAgent sets each simulation round to a fixed real-world duration, typically set to 5 days. (2) Seed Citation Network: An initial citation network, $G_{0}$, is constructed using data from real-world citation networks. Given that graphs are mathematical structures of entity-wise interactions [6], $G_{0}$ encompasses two entity types: author and paper. We collect seven text-rich attributes for papers and five for authors. The authors and papers in $G_{0}$ form the initial author set AA and paper collection set $P$, respectively. (3) Paper Hyperparameter: We set three parameters for paper creation in each simulation round: the number of new papers to be generated, denoted as $n_{p}$; the number of recommended authors for each paper, denoted as $n_{a}$; and the number of recommended citations for each paper, denoted as $n_{k}$.
After setup configuration with $G_{0}$, CiteAgent conducts iterative simulation steps, each simulation round typically representing a five-day period. As shown in Figure 1, each simulation step integrates new papers into $P$ and new authors into AA, following three main stages: Initialization, Socialization, and Creation.
(1) Initialization Stage: This stage involves updating author set AA with heterogeneous LLM-based agents. To achieve a believable simulation of human reference behaviors, we adhere to a general paradigm [32] for constructing LLM-based agents. Having initialized set AA with authors from $G_{0}$, we aim to further expand the quantity and diversity of authors in CiteAgent. We update AA by incorporating new authors and pruning overloaded ones, with the adjustments guided by the authors’ activity frequency. Typically, we add 20 new authors every 5 days, and we designate an author as overloaded if his published papers reach the threshold (typically 100). The newly added authors are initialized with synthetic author profiles generated by LLMs, including nationality, expertise, institution, etc. Each author is designed to emulate human-like behaviors such as thinking, socializing, and drafting, and each author possesses a memory component to store their action trajectory. Detailed prompt templates used at each stage of this process are provided in Section 1 of the supplementary material.
@@
If $|A_{i}^{c}|<n_{a}$, we proceed to the second step. In this step, we filter $A_{i}^{H}$ based on the author profile of $a_{i}$. The system pre-defines a list of author attributes for CRA, denoted as $PF^{c}$. In our experiment, $PF^{c}$ is arranged in the following order: Topic $\rightarrow$ Expertise $\rightarrow$ Citation $\rightarrow$ Institution $\rightarrow$ Nationality. For each attribute $pf\in PF^{c}$, we filter $a_{j}$ from AA if $a_{j}$ shares the same value as $a_{i}$ for the attribute pfpf. Through multiple iterations of this process, we gradually add collaborators to $A_{i}^{c}$.
@@
Next, To simulate the referencing paper search process in real-world academic environments, we develop a scholarly search engine. When first author $a_{i}$ conducts a search to retrieve the most relevant papers, he first generate a query set $Q_{i}$, which is conducted with a list of query words about paper keywords and topics. To get recommended papers $P_{i}^{c}$ based on query words, the process is divided into two steps as outlined in Algorithm 2: recall and rerank. Firstly, For each query word $q\in Q_{i}$, in the recall stage, we filter the top papers using a vector retriever. We begin by constructing a vector database from the embedding vectors of existing papers $P$. Using the SBERT [37] as our embedding model, we embed each paper $p\in P$ into a 384-dimensional vector $e_{p}$. we embed $q$ as a vector $e_{q}$ and calculate the cosine similarity between $e_{q}$ and $e_{p}$ to identify the most relevant papers. This process recommends the top 20 most relevant paper set for each $q$, denoted as $o_{i}^{q}$. Then, in the rerank stage, we reorder $o_{i}^{q}$based on the personal preference of authors, producing a final candidate paper set $P_{i}^{c}$. Similar to CRA, the system predefines a list of paper attributes for the paper recommendation, denoted as $PF^{p}$. For each attribute $pf\in PF^{p}$, we rearrange the paper order in $o_{i}^{q}$ based on the attribute pfpf. In our experiment, $PF^{p}$ is arranged in the following order: Paper Citation $\rightarrow$ Paper Topic. As a result, $o_{i}^{q}$ is first sorted based on citation counts, and then by paper topics.

### v2

@@ 1 Main
Figure 1: An Illustration of One Simulation Step in the CiteAgent Framework. a, Initialization: LLM-based agents are built as distinct authors, each endowed with distinct attributes, denoted as $A$; b, Socialization: We designate an active author subset, denoted as $A_{a}$. Each active agent $a\in A_{a}$ engages in a group discussion with collaborators and collaboratively develops paper drafts. c, Creation: Each active agent utilizes a scholarly search engine to retrieve relevant papers and finalize the paper drafts with reference selection.
@@
CiteAgent includes two entity types: the author set $A={a_{1},\dots,a_{i},\dots a_{n}}$ and the paper set $P={p_{1},\dots,p_{j},\dots p_{m}}$. We employ LLM-based agents to initiate authors and mimic human behaviors. We model interactions between these entities to trace the evolution of citation networks. We initialize the author set $A$ and paper set $P$ from the citation network dataset $G_{0}$, which is enriched with seven paper attributes (e.g., paper content (abstract), paper timeliness, author name, author country, paper citation, paper topic, and author citation) and five author attributes (e.g., citation, institution, expertise, topic, and nationality). Following the $G_{0}$ setup, we run iterative simulation steps in CiteAgent, as shown in Figure 1. Each step integrates new papers and authors into $A$ and $P$ following three main stages: Initialization, Socialization, and Creation. In the initialization stage, we build LLM-based agents to instantiate authors. We update $A$ with new authors, and adjust for overloaded authors who have published excessively. Moving on to the socialization stage, authors collaborate, sharing insights and develop paper drafts. Finally, in the creation stage, authors finalize drafts by referencing relevant papers and updating $P$ with completed papers. As simulation steps proceed, $G_{0}$ progressively expands into a larger network. These steps represent the evolutionary process of the citation network. For validating the simulation authenticity of CiteAgent, we verify that the generated citation network conforms to several classic network science theories, including the power-law distribution [1], citational distortion [22] and shrinking diameter [28] phenomenon.
@@ 2.4 Reasons Behind Citational Distortion
To address this issue, we propose a new metric that normalizes for the cumulative advantage of paper volume. Inspired by the Institutional Preference Score (IPS), which was introduced to capture the citation preferences of various institutions for AI papers in the industry [19]. Likewise, we propose the Referencing Preference Score (RPS) to examine the referencing preferences of agent set $A$ for papers from country set $C$:
@@
Here, $\text{RPS}(C)>1$ demonstrates a preference for papers from $C$ than would be expected under random referencing behaviors. If reference selection is determined by preferential attachment based on paper counts, we would expect the referencing probability to mirror the distribution of these counts, resulting in an RPS of 1 for all countries (if the ref. share of $C$ in papers is zero, we default to $\text{RPS}(C)=1$. We repeat all simulation experiments three times to calculate the RPS standard deviation error. As shown in Figure 5c, for peripheral countries, we find that $\text{RPS}(C_{\text{core}})$ is comparable to, or even slightly lower than, $\text{RPS}(C_{\text{peripheral}})$. Moreover, as illustrated in Figure 5d, in the simulations of GPT-4o-mini and LLAMA-3-70B based authors, $\text{RPS}(C_{\text{peripheral+core}})$ gradually exceeds $\text{RPS}(C_{\text{core}})$ over the years. In contrast, in the simulations of GPT-3.5 based authors, $\text{RPS}(C_{\text{peripheral+core}})$ consistently remains slightly lower than $\text{RPS}(C_{\text{core}})$. Thus, in the CiteAgent simulations, authors exhibit no significant preference for papers from peripheral countries over those from core countries; in fact, they show a slight inclination toward papers from peripheral countries. This observation is consistent with the citation motivation statistics presented in Figure 4a and the results of anonymous-public counterfactual experiments, further confirming the absence of preference for papers from core countries over those from peripheral countries. Therefore, the observed inequality in international citations primarily stems from the uneven distribution of researchers across countries, particularly the concentration of authors in core countries, rather than from intentional biases based on the authors’ nationality during the reference selection.
@@ 4 Methods
Before simulation steps begin, CiteAgent undertakes three configuration setups: (1) Time Flow: CiteAgent sets each simulation round to a fixed real-world duration, typically set to 5 days. (2) Seed Citation Network: An initial citation network, $G_{0}$, is constructed using data from real-world citation networks. Given that graphs are mathematical structures of entity-wise interactions [6], $G_{0}$ encompasses two entity types: author and paper. We collect seven text-rich attributes for papers and five for authors. The authors and papers in $G_{0}$ form the initial author set $A$ and paper collection set $P$, respectively. (3) Paper Hyperparameter: We set three parameters for paper creation in each simulation round: the number of new papers to be generated, denoted as $n_{p}$; the number of recommended authors for each paper, denoted as $n_{a}$; and the number of recommended citations for each paper, denoted as $n_{k}$.
After setup configuration with $G_{0}$, CiteAgent conducts iterative simulation steps, each simulation round typically representing a five-day period. As shown in Figure 1, each simulation step integrates new papers into $P$ and new authors into $A$, following three main stages: Initialization, Socialization, and Creation.
(1) Initialization Stage: This stage involves updating author set $A$ with heterogeneous LLM-based agents. To achieve a believable simulation of human reference behaviors, we adhere to a general paradigm [32] for constructing LLM-based agents. Having initialized set $A$ with authors from $G_{0}$, we aim to further expand the quantity and diversity of authors in CiteAgent. We update $A$ by incorporating new authors and pruning overloaded ones, with the adjustments guided by the authors’ activity frequency. Typically, we add 20 new authors every 5 days, and we designate an author as overloaded if his published papers reach the threshold (typically 100). The newly added authors are initialized with synthetic author profiles generated by LLMs, including nationality, expertise, institution, etc. Each author is designed to emulate human-like behaviors such as thinking, socializing, and drafting, and each author possesses a memory component to store their action trajectory. Detailed prompt templates used at each stage of this process are provided in Section 1 of the supplementary material.
@@
If $|A_{i}^{c}|<n_{a}$, we proceed to the second step. In this step, we filter $A_{i}^{H}$ based on the author profile of $a_{i}$. The system pre-defines a list of author attributes for CRA, denoted as $PF^{c}$. In our experiment, $PF^{c}$ is arranged in the following order: Topic $\rightarrow$ Expertise $\rightarrow$ Citation $\rightarrow$ Institution $\rightarrow$ Nationality. For each attribute $pf\in PF^{c}$, we filter $a_{j}$ from $A$ if $a_{j}$ shares the same value as $a_{i}$ for the attribute $pf$. Through multiple iterations of this process, we gradually add collaborators to $A_{i}^{c}$.
@@
Next, To simulate the referencing paper search process in real-world academic environments, we develop a scholarly search engine. When first author $a_{i}$ conducts a search to retrieve the most relevant papers, he first generate a query set $Q_{i}$, which is conducted with a list of query words about paper keywords and topics. To get recommended papers $P_{i}^{c}$ based on query words, the process is divided into two steps as outlined in Algorithm 2: recall and rerank. Firstly, For each query word $q\in Q_{i}$, in the recall stage, we filter the top papers using a vector retriever. We begin by constructing a vector database from the embedding vectors of existing papers $P$. Using the SBERT [37] as our embedding model, we embed each paper $p\in P$ into a 384-dimensional vector $e_{p}$. we embed $q$ as a vector $e_{q}$ and calculate the cosine similarity between $e_{q}$ and $e_{p}$ to identify the most relevant papers. This process recommends the top 20 most relevant paper set for each $q$, denoted as $o_{i}^{q}$. Then, in the rerank stage, we reorder $o_{i}^{q}$based on the personal preference of authors, producing a final candidate paper set $P_{i}^{c}$. Similar to CRA, the system predefines a list of paper attributes for the paper recommendation, denoted as $PF^{p}$. For each attribute $pf\in PF^{p}$, we rearrange the paper order in $o_{i}^{q}$ based on the attribute $pf$. In our experiment, $PF^{p}$ is arranged in the following order: Paper Citation $\rightarrow$ Paper Topic. As a result, $o_{i}^{q}$ is first sorted based on citation counts, and then by paper topics.